The sum of two numbers is the same as four times the smaller number. If twice the larger is decreased by the smaller, the result is 30. Find the numbers.
Answer by stanbon(48575)

The sum of two numbers is the same as four times the smaller number. If twice the larger is decreased by the smaller, the result is 30. Find the numbers
---------------------------------------
Equations:
x + y = 4x
2y-x = 30
-----------------
Rearrange:
y = 3x
x = 2y - 30
substitute to solve for "Y":
y = 3(2y-30)
y = 6y - 90
5y = 90
y = 18
------------
x = 2y-30 = 2*18-30 = 6
------------------------------
Cheers,
StanH.
=================